OpenStudy (anonymous):

Can someone please explain the two solutions for the following equation? x^2-9=0

jimthompson5910 (jim_thompson5910):

9 is a perfect square because 9 = 3^2

jimthompson5910 (jim_thompson5910):

so x^2-9 is really x^2-3^2 notice we have a difference of squares

jimthompson5910 (jim_thompson5910):

So we factor the left side using the difference of squares formula x^2-9=0 x^2-3^2=0 (x-3)(x+3) = 0 what's next?

OpenStudy (anonymous):

you dont get it..?

OpenStudy (precal):

also, if you look at the graph. The solutions for this quadratic is your x intercepts.